# Entanglement in the Schwinger effect

### v1

@@ IV.2.5 Axial anomaly
which is compatible with the axial anomaly [anomaliesQFT] and reproduces the well-known independence of the anomaly on the temperature at equilibrium [Reuter1985, Dolan:1973].
In terms of entanglement, it is interesting to note that it is possible to choose the external parameters such that we only have entanglement in one of the two sectors
$(\omega^{(L)})^{2}=k_{z}^{2}+2(l+1)eB>(\omega^{(R)})^{2}=k_{z}^{2}+2l\,$, (126)
which implies that the maximum temperature to have entanglement will differ
$T_{\text{max}}^{(L)}=\frac{\omega^{(L)}}{\text{arcsinh}(1)}>T_{\text{max}}^{(R)}=\frac{\omega^{(R)}}{\text{arcsinh}(1)}\,.$ (127)
The maximum temperature for the left sector is always $\Delta T=2B/\text{arcsin}(1)$ higher that the maximum temperature for the right sector. For example, the pair with $k_{z}=0$ and $l=1$ in the left sector will be entangled at $T=3eB/\text{arcsin}(1)$ while the pair with the same quantum numbers in the right sector won’t. [^16] See Ref. [Florio:2025xup] for a detailed analysis of an entanglement asymmetry in the massless Schwinger model in 1+1.
IV.2.6 Mutual information
@@ IV.3 Connection to experiment
In vacuum QED, the relevant scale for the Schwinger effect is the Schwinger field $E_{S}=\frac{m_{e}^{2}c^{3}}{e\hbar}\approx 1.3\times 10^{18}\mathrm{\penalty 10000\ V}/\mathrm{m}$, which corresponds to an intensity of $I_{S}\approx 10^{29}\mathrm{\penalty 10000\ W}/\mathrm{cm}^{2}$. Current state-of-the-art lasers can reach intensities up to $I\approx 10^{23}\mathrm{\penalty 10000\ W}/\mathrm{cm}^{2}$ (see, for example Refs. [Turcu:2016dxm, Yoon:2021ony]), i.e., $E\approx 10^{15}\mathrm{\penalty 10000\ V}/\mathrm{m}$, still several orders of magnitude below $E_{S}$. [^17] Various schemes (e.g. pulse shaping or dynamical assistance) have been proposed to enhance strong-field signatures, but even optimistic projections remain below $E_{S}$ (see, e.g., [Schutzhold:2008pz, Bulanov:2010gb, Aleksandrov:2022wcz]). Therefore, a direct observation of vacuum pair creation remains beyond present capabilities. In this regime, it is also not meaningful to discuss pair entanglement: measurements capture total production and current, not correlations of specific pairs. This motivates the search for analogue realizations where the same physics can be probed under experimentally accessible conditions.
@@
In the mesoscopic Schwinger effect, an applied electric field in graphene induces electron-hole pair production, the solid-state analogue of vacuum pair creation. In this context, our formulas directly apply with the replacements $mc^{2}\rightarrow\Delta$ (the effective gap, i.e., the minimum energy scale that appears in the channel due to its finite size) and $c\rightarrow v_{F}$ (the Fermi velocity), with the caveat that the experiment is effectively in $1+2$ rather than $1+3$ dimensions. This dimensionality change does not alter the mode-by-mode analysis we use, but it does affect integrated quantities such as the total current (see Ref. [Kim:2000un] for explicit results in $d$ dimensions). It was recently reported universal one-dimensional Schwinger conductance [Schmitt:2022pkd], consistent with electron–hole pair production in graphene. For the effective gap values quoted in the experiment, namely $\Delta\lesssim 0.2\mathrm{\penalty 10000\ eV}$, the associated Schwinger field is $E_{S}=\Delta^{2}/e\hbar v_{\mathrm{F}}\approx 6\times 10^{7}\mathrm{\penalty 10000\ V}/\mathrm{m}$, which, as discussed, is experimentally accessible. The same experimental setting is also promising for probing entanglement. Focusing on the lowest-energy mode ($k_{\perp}=0$ and $\omega=\Delta$), the maximum temperature allowed for fermionic entanglement is of the order of $T_{\max}\approx 2500\mathrm{\penalty 10000\ K}$, so room temperature ($300\mathrm{\penalty 10000\ K}$) lies far below this threshold. In that case, the entanglement window is broad: for $\Delta=0.2\mathrm{eV}$ at $T=300\mathrm{\penalty 10000\ K}$ one finds $E_{\min}\approx 1.2\times 10^{7}\mathrm{\penalty 10000\ V}/\mathrm{m}$ and $E_{\max}\approx 10^{15}\mathrm{\penalty 10000\ V}/\mathrm{m}$, centered at the optimal value for entanglement $E_{\star}\approx 2.7\times 10^{8}\mathrm{\penalty 10000\ V}/\mathrm{m}$ [see eqs. (120) and (121)]. By contrast, if the gap were much smaller, for example $\Delta=20\,\mathrm{\penalty 10000\ meV}$, then $T_{\max}\approx 200\mathrm{\penalty 10000\ K}$, so at room temperature entanglement wouldn’t be possible, although pairs could still be produced.
@@
Finally, we would like to briefly mention an experimental possibility to detect the analogue Schwinger effect for bosons. In Refs. [Hongo:2020xaw, Adorno:2023olb] it has been proposed that magnetic inhomogeneities in chiral magnets can trigger magnon–antimagnon pair production in close analogy to electron–positron creation in vacuum. In this mapping, the role of the electric field is played by the external magnetic profile, while magnons take the place of charged particles (i.e., the formulas above for pure electric field $E$ can be used with the replacements $eE\to\mu_{\text{eff}}\nabla B$ with $\mu_{\text{eff}}$ the effective magnetic moment of a magnon, $m_{e}c^{2}\to\Delta$ with $\Delta$ the magnon gap set by the experiment, and $c\to v_{m}$ with $v_{m}$ the spin–wave velocity). In this setting, the protocol we have outlined in subsection IV.1.5 to maximize entanglement through squeezed initial states could also be implemented. In fact, proposals to engineer squeezed magnon states already exist in related cavity platforms [Li:2019cxy], suggesting that this route might provide a realistic option to test quantum correlations in a bosonic analog of the Schwinger mechanism.
@@ V Conclusions
For fermions, Pauli blocking competes with pair creation, producing the characteristic $\tanh(\omega/2T)$ in the particle number (see eq. (105)). Entanglement exhibits a qualitatively different pattern with respect to the scalar case: at a fixed $T$, and provided $T<T_{\text{max}}=\omega/\text{arcsinh}(1)$, a two-mode pair will be entangled only if the external field is inside a finite window $E_{\text{min}}<E<E_{\text{max}}$, with the limiting values given in eq. (120). There is an optimal value of the electric field for entanglement $E=E_{*}$ (see eq. (121)), which corresponds to the value that maximizes the logarithmic negativity. As a consistency check, we reproduced the axial anomaly and commented on the entanglement differences between the two sectors ($L$ and $R$).
@@ Acknowledgements.
AM and SP thank Björn Garbrecht for useful comments. The work of SP was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2094 – 390783311. DK acknowledges financial support by DIM ORIGINES program from Région Île de France. DK also received support through the Atracción de Talento Cesar Nombela grant No 2023-T1/TEC-29023, funded by Comunidad de Madrid (Spain); as well as financial support via the Spanish Grant PID2023-149560NB-C21, funded by MCIU/AEI/10.13039/501100011033/FEDER, UE.
@@ A.2 Vacuum persistence amplitude from the Bogoliubov coefficients
Since we work with fermions, the average number of particles equals the probability of producing a pair $\langle N_{I}\rangle=|\beta_{I}|^{2}=P(1)$. Therefore we can compute the vacuum-to-vacuum probability as P(0)=1P(1)P(0)=1-P(1) for each mode (for more details see, for example, Ref. [parker-toms], chapter 2). Therefore, the vacuum persistence amplitude (taking into account all modes) reads
@@
where we used the changes of variables $\text{d}k_{y}=-eB\text{d}x$ and $\text{d}k_{z}=-eE\text{d}t$ and thus $\int\text{d}k_{y}=|eB|L$ and $\int\text{d}k_{z}=|eE|\Delta t$ where $L=V^{1/3}$. This result is in agreement with the standard formula [Popov:1971iga, Dunne:2004nc].
@@ Appendix B Gaussian formalism for fermions/details for fermionic systems
Although the fermionic Gaussian formalism is not needed for the observables discussed in this work, since all relevant quantities can be directly computed from the finite density matrix (see Section III.3), it provides a compact and parallel way to treat bosons and fermions in the same language. We will give here an operational approach. For a formal discussion, see, for example, Ref. [Hackl:2020ken]. As in the scalar case, for a pair of fermionic modes, their respective ladder operators $\{\hat{b}_{1},\hat{b}^{\dagger}_{1},\hat{b}_{2},\hat{b}^{\dagger}_{2}\}$ can be arranged in a 4-vector $\bm{\psi}_{J}$. After interaction, in the Heisenberg picture, the $^{(\text{in})}$ and $^{(\text{out})}$ operators are related by a Bogoliubov transformation. In equations
@@
We do not use the Gaussian formalism to compute the logarithmic negativity. For bosons, partial transposition (PT) is a linear map that preserves gaussianity. For fermions, PT does not preserve gaussianity (because of anticommutation), so it is more direct to compute the logarithmic negativity directly from the density matrix, as we did in the main text.

### v2

@@ IV.2.6 Axial anomaly
which is compatible with the axial anomaly [13] and reproduces the well-known independence of the anomaly on the temperature at equilibrium [64, 20].
In terms of entanglement, it is interesting to note that the fermionic logarithmic negativity exhibits different magnitudes in the two chiral sectors because their effective frequencies differ by one Landau level. In particular,
$(\omega^{(L)}_{l})^{2}=k_{z}^{2}+m^{2}+2(l+1)eB\;>\;(\omega^{(R)}_{l})^{2}=k_{z}^{2}+m^{2}+2leB\,.$ (132)
More concretely, the fermionic logarithmic negativity satisfies $L_{N}^{(L)}(l)=L_{N}^{(R)}(l+1)$ mode by mode. Therefore, when comparing the two sectors at fixed Landau level $l$, $L_{N}$ can be parametrically different in the $L$ and $R$ sectors for the same external parameters. Operationally, if one imposes a finite experimental threshold $L_{N}>\ln(1+\epsilon^{2})/\ln 2$, it is possible to have entanglement above threshold in one chiral sector while it lies below threshold in the other. [^17] See Ref. [29] for a detailed analysis of an entanglement asymmetry in the massless Schwinger model in 1+1.
IV.2.7 Mutual information
@@ IV.3 Connection to experiment
In vacuum QED, the relevant scale for the Schwinger effect is the Schwinger field $E_{S}=\frac{m_{e}^{2}c^{3}}{e\hbar}\approx 1.3\times 10^{18}\mathrm{\penalty 10000\ V}/\mathrm{m}$, which corresponds to an intensity of $I_{S}\approx 10^{29}\mathrm{\penalty 10000\ W}/\mathrm{cm}^{2}$. Current state-of-the-art lasers can reach intensities up to $I\approx 10^{23}\mathrm{\penalty 10000\ W}/\mathrm{cm}^{2}$ (see, for example Refs. [76, 80]), i.e., $E\approx 10^{15}\mathrm{\penalty 10000\ V}/\mathrm{m}$, still several orders of magnitude below $E_{S}$. [^18] Various schemes (e.g. pulse shaping or dynamical assistance) have been proposed to enhance strong-field signatures, but even optimistic projections remain below $E_{S}$ (see, e.g., [67, 18, 7]). Therefore, a direct observation of vacuum pair creation remains beyond present capabilities. In this regime, it is also not meaningful to discuss pair entanglement: measurements capture total production and current, not correlations of specific pairs. This motivates the search for analogue realizations where the same physics can be probed under experimentally accessible conditions.
@@
In the mesoscopic Schwinger effect, an applied electric field in graphene induces electron-hole pair production, the solid-state analogue of vacuum pair creation. In this context, our formulas directly apply with the replacements $mc^{2}\rightarrow\Delta$ (the effective gap, i.e., the minimum energy scale that appears in the channel due to its finite size) and $c\rightarrow v_{F}$ (the Fermi velocity), with the caveat that the experiment is effectively in $1+2$ rather than $1+3$ dimensions. This dimensionality change does not alter the mode-by-mode analysis we use, but it does affect integrated quantities such as the total current (see Ref. [47] for explicit results in $d$ dimensions). It was recently reported universal one-dimensional Schwinger conductance [65], consistent with electron–hole pair production in graphene. For the effective gap values quoted in the experiment, namely $\Delta\lesssim 0.2\mathrm{\penalty 10000\ eV}$, the associated Schwinger field is $E_{S}=\Delta^{2}/e\hbar v_{\mathrm{F}}\approx 6\times 10^{7}\mathrm{\penalty 10000\ V}/\mathrm{m}$, which, as discussed, is experimentally accessible. The same experimental setting is also promising for probing entanglement. Focusing on the lowest-energy mode ($k_{\perp}=0$ and $\omega=\Delta$), the fermionic logarithmic negativity does not exhibit a sharp critical temperature, but it is smoothly suppressed as $T/\omega$ increases. In practice, entanglement detection is limited by experimental resolution, so it is natural to adopt an operational criterion $L_{N}>L_{N}^{\text{min}}(\epsilon)$ as discussed around (123) that will give an operational window for entanglement detection $E_{-}(\epsilon)<E<E_{+}(\epsilon)$ centered around the optimal value $E_{\star}\approx 2.7\times 10^{8}\mathrm{\penalty 10000\ V}/\mathrm{m}$. At room temperature $T=300\mathrm{\penalty 10000\ K}$ and for $\epsilon=10^{-2}$, one finds $E_{-}\approx 1.7\times 10^{7}\mathrm{\penalty 10000\ V}/\mathrm{m}$ and $E_{+}\approx 1.26\times 10^{13}\mathrm{\penalty 10000\ V}/\mathrm{m}$ so that $E_{S}$ would fit within the right window.
@@
Finally, we would like to briefly mention an experimental possibility to detect the analogue Schwinger effect for bosons. In Refs. [42, 2] it has been proposed that magnetic inhomogeneities in chiral magnets can trigger magnon–antimagnon pair production in close analogy to electron–positron creation in vacuum. In this mapping, the role of the electric field is played by the external magnetic profile, while magnons take the place of charged particles (i.e., the formulas above for pure electric field $E$ can be used with the replacements $eE\to\mu_{\text{eff}}\nabla B$ with $\mu_{\text{eff}}$ the effective magnetic moment of a magnon, $m_{e}c^{2}\to\Delta$ with $\Delta$ the magnon gap set by the experiment, and $c\to v_{m}$ with $v_{m}$ the spin–wave velocity). In this setting, the protocol we have outlined in subsection IV.1.5 to maximize entanglement through squeezed initial states could also be implemented. In fact, proposals to engineer squeezed magnon states already exist in related cavity platforms [48], suggesting that this route might provide a realistic option to test quantum correlations in a bosonic analog of the Schwinger mechanism.
@@ V Conclusions
For fermions, Pauli blocking competes with pair creation, producing the characteristic $\tanh(\omega/2T)$ in the particle number (see eq. (106)). Entanglement exhibits a qualitatively different pattern with respect to the scalar case: the logarithmic negativity is a smooth function of $T$ and $|\beta_{I}|^{2}$ (which is constrained between 0 and 1). It is monotonically suppressed as $T$ increases, and, at fixed $T$, it reaches its maximum at $\left|\beta_{I}\right|^{2}=1/2$, which defines a temperature-independent optimal field $E=E_{*}$ (see eq. (122)). For experimental purposes, one may nonetheless introduce an operational thereshold $L_{N}>\epsilon$, which induces an effective entanglement window $E_{-}(\epsilon)<E<E_{+}(\epsilon)$ and a corresponding temperature scale $T_{c}(\epsilon)$ (see eq. (123) and the subsequent discussion). As a consistency check, we reproduced the axial anomaly and commented on the entanglement differences between the two sectors ($L$ and $R$).
@@ Acknowledgements.
The authors thank Björn Garbrecht, Pablo Chisvert, and Ivan Agullo for their insightful comments. The work of SP was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2094 – 390783311. DK acknowledges financial support by DIM ORIGINES program from Région Île de France. DK also received support through the Atracción de Talento Cesar Nombela grant No 2023-T1/TEC-29023, funded by Comunidad de Madrid (Spain); as well as financial support via the Spanish Grant PID2023-149560NB-C21, funded by MCIU/AEI/10.13039/501100011033/FEDER, UE.
@@ A.2 Vacuum persistence amplitude from the Bogoliubov coefficients
Since we work with fermions, the average number of particles equals the probability of producing a pair $\langle N_{I}\rangle=|\beta_{I}|^{2}=P(1)$. Therefore we can compute the vacuum-to-vacuum probability as P(0)=1P(1)P(0)=1-P(1) for each mode (for more details see, for example, Ref. [56], chapter 2). Therefore, the vacuum persistence amplitude (taking into account all modes) reads
@@
where we used the changes of variables $\text{d}k_{y}=-eB\text{d}x$ and $\text{d}k_{z}=-eE\text{d}t$ and thus $\int\text{d}k_{y}=|eB|L$ and $\int\text{d}k_{z}=|eE|\Delta t$ where $L=V^{1/3}$. This result is in agreement with the standard formula [62, 23].
@@ Appendix B Gaussian formalism for fermions/details for fermionic systems
Although the fermionic Gaussian formalism is not needed for the observables discussed in this work, since all relevant quantities can be directly computed from the finite density matrix (see Section III.3), it provides a compact and parallel way to treat bosons and fermions in the same language. We will give here an operational approach. For a formal discussion, see, for example, Ref. [37]. As in the scalar case, for a pair of fermionic modes, their respective ladder operators $\{\hat{b}_{1},\hat{b}^{\dagger}_{1},\hat{b}_{2},\hat{b}^{\dagger}_{2}\}$ can be arranged in a 4-vector $\bm{\psi}_{J}$. After interaction, in the Heisenberg picture, the $^{(\text{in})}$ and $^{(\text{out})}$ operators are related by a Bogoliubov transformation. In equations
@@
We note that for physical states $\bm{\mu}=0$ as discussed, for example, in Ref. [70].
@@
Appendix C Operational lower bound in Logarithmic Negativity
In this appendix, we derive the error in determining the logarithmic negativity given the experimental error $\delta$ in the construction of the density matrix.
Consider a $d\times d$ density matrix $\hat{\rho}$ with $s\leq d^{2}$ non-zero entries. Let $\hat{\rho}^{\prime}$ be the reconstructed density matrix from an experiment that is supposed to match $\hat{\rho}$. Let $\delta_{ij}$ be the error bars of the (theoretically) expected density matrix elements, i.e.
$|\hat{\rho}_{ij}-\hat{\rho}_{ij}^{\prime}|\leq\delta_{ij},\quad\text{for}\,(i,j)\in S.$ (154)
where $S=\{(i,j)\in\{1,...,d\}^{2}:\hat{\rho}_{ij}\neq 0\}$, and let $s\leq d^{2}$ be the number of the non-zero elements of $\hat{\rho}$ (and $\hat{\rho}^{\prime}$).
The logarithmic negativity $L_{N}$, measuring the entanglement across a partition in the state $\hat{\rho}$, is defined as
$L_{N}(\rho)=\log_{2}\!\left\|\hat{\rho}^{\text{PT}}\right\|_{1}\,.$ (155)
Here, $\hat{\rho}^{\text{PT}}$ represents the partially transposed density matrix for bosons, and the partially time reversed density matrix for fermions. $\hat{X}||_{1}:=\text{Tr}\sqrt{\hat{X}^{\dagger}\hat{X}}$ is the trace norm of the operator $\hat{X}$.
We are interested in the error in logarithmic negativity, i.e. we want to find an upper bound in
[…]
Notice that from (154) and assuming, for simplicity, a uniform error across the non-zero entries, $\delta_{ij}=\delta$, we can write
$|\hat{\rho}^{\text{PT}}_{ij}-\hat{\rho}_{ij}^{{}^{\prime}\text{PT}}|\leq\delta\,.$ (157)
To start, it is easier to compute the Frobenius norm of the partially transposed density matrix difference and then relate this to the corresponding trace norms. The Frobenius norm of an operator $\hat{X}$ with matrix entries $X_{ij}$ is
$\hat{X}||_{2}:=\sqrt{\sum_{i,j}|X_{ij}|^{2}}.$ (158)
Setting $\hat{X}=\hat{\rho}^{\text{PT}}-\hat{\rho}^{{}^{\prime}\text{PT}}$
$\hat{\rho}^{\text{PT}}-\hat{\rho}^{{}^{\prime}\text{PT}}||_{2}:=\sqrt{\sum_{i,j}|\hat{\rho}_{ij}^{\text{PT}}-\hat{\rho}^{{}^{\prime}\text{PT}}_{ij}|^{2}}\leq\sqrt{s}\delta,$ (159)
where we made use of the fact that $\hat{\rho}$ and $\hat{\rho}^{{}^{\prime}\text{PT}}$ have $s$ non-zero entries.
Now, we want to connect the Frobenius to the trace norm since it is the latter that enters in the definition of logarithmic negativity (155). To this end, we use the inequality (3) of reference [19]
[…]
For $\hat{X}=\hat{\rho}^{\text{PT}}-\hat{\rho}^{{}^{\prime}\text{PT}}$, with $\text{rank}(\hat{\rho}^{\text{PT}}-\hat{\rho}^{{}^{\prime}\text{PT}})\leq d$, we get
[…]
Thus, from (159) and (161), one finds
$\hat{\rho}^{\text{PT}}-\hat{\rho}^{{}^{\prime}\text{PT}}||_{1}\leq\sqrt{sd}\,\delta$ (162)
Furthermore, the several norms obey the triangle inequality. For instance, for the trace norm of two operators $\hat{X},\hat{Y}$, we have
[…]
Setting $\hat{X}=\hat{\rho}^{\text{PT}}$ and $\hat{Y}=\hat{\rho}^{{}^{\prime}\text{PT}}-\hat{\rho}^{\text{PT}}$, we find
[…]
Similarly, setting $\hat{X}=\hat{\rho}^{{}^{\prime}\text{PT}}$ and $\hat{Y}=\hat{\rho}^{\text{PT}}-\hat{\rho}^{{}^{\prime}\text{PT}}$, we find
[…]
Combining the last two inequalities and applying (162), one finds
[…]
Furthermore, one can show, e.g. using the mean value theorem, that for two real numbers $a$, $b$, satisfying $1\leq a\leq b$,
$\ln b-\ln a\leq b-a,$ (167)
[…]
Finally, combining (166) with (168), we obtain
$\left|L_{N}(\hat{\rho})-L_{N}(\hat{\rho}^{\prime})\right|\leq\frac{\sqrt{sd}\delta}{\ln 2}.$ (169)
We define the threshold parameter
$\epsilon(\delta):=\frac{\sqrt{sd}}{\ln 2}\delta.$ (170)
[…]
$L_{N}(\hat{\rho})>\epsilon(\delta)$ (171)
to guarantee entanglement detection, i.e. $L_{N}(\hat{\rho}^{\prime})>0$. Hence, $\epsilon(\delta)$ is the threshold value for $L_{N}$.
For the out density matrix in the fermionic Schwinger effect (103), there are only $6$ non-zero elements of $\hat{\rho}$. Plugging $d=4$, $s=6$, we obtain
$L_{N}(\hat{\rho})\gtrsim\frac{2\sqrt{6}}{\ln 2}\delta=7.07\delta.$ (172)